USE THE METHOD OF DIRECT TO PROVE:


If a is an odd integer, then a’2 + 3a +5 is odd.


1
Expert's answer
2022-02-14T16:30:57-0500

Thus "a = 2c +1" for some integer "c," by definition of an odd number. Then


"a^2+3a+5=(2c+1)^2+3(2c+1)+5"

"=4c^2+4c+1+6c+3+5=4c^2+10c+9"

"=2(2c^2+5c+4)+1=2b+1,"

where "b=2c^2+5c+4."

Therefore by definition of an odd number, if "a" is an odd integer, then "a^2 + 3a +5" is odd.